Berman & Jaccor Corporation's current sales and partial balance sheet are shown below. 


This year 
Sales$1,000
Balance Sheet: Assets Cash$ 150 
 Short-term investments  $ 145
Accounts receivable  $ 300
Inventories  $ 300
Total current assets $ 895
Net fixed assets  $ 600
Total assets  $ 1,495 

Sales are expected to grow by 12% next year. Assuming no change in operations from this year to next year, what are the projected total operating assets? Do not round intermediate calculations. Round your answer to the nearest dollar.
Answer #1

Operating assets are cash, accounts receivable, inventories and net fixed assets.

First we calculate ratio of each operating assets of last year to last year's sales.

Cash = (cash/sales) = $150/$1,000 = 0.15

Accounts receivable = $300/$1,000 = 0.3

Inventories = $300/$1,000 = 0.3

Net fixed assets = $600/$1,000 = 0.60

Next year's sales = current sales*(1+growth rate) = $1,000*1.12 = $1,120

Project cash next year = project sales*cash to sales ratio = $1,120*0.15 = $168

Project accounts receivable = $1,120*0.3 = $336

Projected inventories = $1,120*0.3 = $336

Projected Net fixed assets = $1,120*0.60 = $672

projected total operating assets = $168 + $336 + $336 + $672 = $1,512